Suppose X is normally distributed with mean 4 and standard deviation 4. Find the probability that 2X exceeds 7.

Group of answer choices

.5497

Suppose a random sample of 25 measurements is taken from a population with mean 17 and variance 100. The variance of the sample mean is

Group of answer choices

2

0.68

4

17

100

.301

.4012

.4555

.5988
Answer #1

Solution :

Given that ,

mean = = 4

standard deviation = = 4

P(x > 7) = 1 - P(x < 7)

= 1 - P[(x - ) / < (7 - 4) / 4)

= 1 - P(z < 0.75)

= 1 - 0.7734

= 0.2266

Probability = 0.2266

mean = = 17

standard deviation = = 10

n = 25

= 17

= / n = 10 / 25 = 2